# Greedy Algorithm for Structured Bandits: A Sharp Characterization of Asymptotic Success / Failure*

### v1

@@ 3 Characterization for multi-armed bandits
It is easy to see that the two definitions are equivalent, in the following sense:
@@
The complexity of the problem instance $(f^{*},\mathcal{F})$ enters our results via its function-gap,
@@
Remark 3.4.
Thus, $\mathtt{Greedy}$ succeeds, in the sense of achieving sublinear regret for any warm-up data, if and only if the problem instance is self-identifiable. Else, $\mathtt{Greedy}$ fails if the warm-up data consists of one sample per arm, as per Theorem 3.3(b), incurring linear expected regret: $\operatornamewithlimits{\mathbb{E}}\left[\,R(t)\,\right]\geq(t-T_{0})\cdot p_{\mathtt{dec}}\cdot(f^{*}(a^{*})-f^{*}(a_{\mathtt{dec}}))$ for each round $t\in(T_{0},\infty)$, where $a^{*}$ is the best arm.
Remark 3.5.
$\mathtt{Greedy}$ succeeds because self-identifiability makes the problem intrinsically “easy.” More precisely, consider any bandit algorithm that avoids playing an arm once it is identified, with high confidence, as both suboptimal and having a specific expected reward. This defines a mild yet fundamental non-degeneracy condition: a reasonable algorithm should never take an action that provides neither new information (exploration) nor utility from existing information (exploitation), whether it prioritizes one or balances both. The class of algorithms satisfying this condition is broad—for instance, an algorithm may continue playing some arm $a$ indefinitely as long as the reward structure permits this arm with expected reward $f^{*}(a)$ to be optimal. However, under self-identifiability, any algorithm satisfying this condition achieves sublinear regret (see Section 7 for formal results and discussion).
The failure probability in Theorem 3.3(b) could be quite low. As a (very) partial mitigation, we observe that when there are multiple decoys $f_{\mathtt{dec}}\in\mathcal{F}$, we could pick one (in the analysis) which maximizes function-gap $\Gamma(f_{\mathtt{dec}})$.
Remark 3.6.
For the sake of a stylized example which admits a more efficient analysis, suppose the decoy arm has deterministic rewards (and all other arms have unit-variance Gaussian rewards, as before). Then, as we show in Appendix D, the failure probability satisfies $p_{\mathtt{dec}}\geq\Omega\left(\,(\Delta(f_{\mathtt{dec}})/\sqrt{K})^{K-1}\,\right)$, where $\Delta(f)$ is the “standard” notion of gap in MAB: the difference between the best and second-best expected reward.
Proof Sketch for Theorem 3.3(a). We show that a suboptimal arm $a$ cannot be chosen more than $\widetilde{O}(K/\Gamma^{2}(f^{*}))$ times throughout the main stage. Indeed, suppose $a$ is chosen this many times by some round $t>T_{0}$. Then $\bar{r}_{t}(a)$, the empirical mean reward for $a$, is within $\Gamma(f^{*})/2$ of its true mean $f^{*}(a)$ with high probability, by a standard concentration inequality. This uniquely identifies $f^{*}(a)$ by definition of the function-gap, which in turn identifies $a$ as a suboptimal arm for any feasible reward function. Intuitively, this should imply that $a$ cannot be chosen again. Making this implication formal requires an additional argument invoking $\mathtt{MSE}_{t}(\cdot)$, as defined in (2.3).
@@
Proof Sketch for Theorem 3.3(b). We are interested in two events. The first event ensures that the warm-up data misidentifies $f_{\mathtt{dec}}$ as the true reward function for all non-decoy arms:
@@
Recall that this is a statement about the single warm-up sample for each non-decoy arm.
The second event asserts that the decoy arm $a_{\mathtt{dec}}$ has empirical mean rewards that are tightly concentrated around their mean, $f^{*}(a_{\mathtt{dec}})$, for all rounds after the warm-up. Specifically:
@@
We note that under $E_{1}\cap E_{2}$, $\mathtt{Greedy}$ always chooses the decoy arm. We lower-bound $\Pr\left[\,E_{1}\cap E_{2}\,\right]$ by observing that the two events are independent (as they concern, resp., $a_{\mathtt{dec}}$ and all other arms) and analyzing each event separately.
4 Characterization for contextual bandits
This section considers $\mathtt{StructuredCB}$: contextual bandits with a known reward structure $\mathcal{F}$. The ideas from Section 3 need to be modified to accommodate contexts. In particular, the trivial reduction to bandits — treat each contexts-to-arms mapping as a “super-arm” in a bandit problem and reduce to $\mathtt{StructuredMAB}$— does not work because, in contrast to the MAB case, $\mathtt{Greedy}$ now observes contexts. Further, such reduction would replace the dependence on $K$ in Theorem 3.3 with the number of mappings, i.e., $K^{|\mathcal{X}|}$, whereas we effectively replace it with $K\cdot|\mathcal{X}|$.
Let’s start with some additional notation for CB. Mappings from contexts to arms are commonly called policies. Let $\Pi$ denote the set of all policies. Expected reward of policy $\pi\in\Pi$ is $f^{*}(\pi):=\operatornamewithlimits{\mathbb{E}}_{x}\left[\,f^{*}(x,\pi(x))\,\right]$, where the expectation is over the fixed distribution of context arrivals. A policy $\pi$ is called optimal for reward function $f$ if it maximizes expected reward $f(\cdot)$, and suboptimal otherwise. Note that an optimal policy satisfies $\pi(x)\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}f(x,\cdot)$ for each context $x$, which is unique by assumption. Let $\pi^{*}$ denote the optimal policy for $f^{*}$.
We work with an alternative description of $\mathtt{Greedy}$ in terms of policies: it chooses policy $\pi_{t}$ in each round $t$, before seeing the context $x_{t}$, and then chooses arm $a_{t}=\pi_{t}(x_{t})$. Here $\pi_{t}$ is the optimal policy for the reward function $f_{t}$ from Eq. (2.1).
@@ 4 Characterization for contextual bandits
Fix a problem instance $(f^{*},\mathcal{F})$ of $\mathtt{StructuredCB}$.
@@
    Suppose the warm-up data consists of one sample for each context-arm pair. Assume $f^{*}$ has a decoy $f_{\mathtt{dec}}\in\mathcal{F}$, with decoy policy $\pi_{\mathtt{dec}}$. Then with some probability $p_{\mathtt{dec}}>0$, $\mathtt{Greedy}$ chooses $\pi_{\mathtt{dec}}$ in all rounds $t\in(T_{0},\infty)$. We have $p_{\mathtt{dec}}\geq X^{-O(KX/\Gamma^{2}(f_{\mathtt{dec}}))}$, where $X=|\mathcal{X}|$.
@@
As in Section 3, $\mathtt{Greedy}$ succeeds (i.e., achieves sublinear regret for any warm-up data) if and only if the problem instance is self-identifiable. Else, $\mathtt{Greedy}$ fails if the warm-up data consists of one sample per each context-arm pair, as per Theorem 4.3(b), incurring linear regret:
@@
Self-identifiability makes the problem “easy”, as in Remark 3.5 (see Section 7).
Proof Sketch for Theorem 4.3(a). Let $X=|\mathcal{X}|$. Directly applying the proof technique from the MAB case results in a regret bound that is linear in $|\Pi|=K^{X}$. Instead, we apply a potential argument and achieve regret bound that is polynomial in $KX$. First, by a standard concentration inequality, we argue that the loss for $f^{*}$ can be upper-bounded as $\mathtt{MSE}_{t}(f^{*})\leq\widetilde{O}(KX)$ with high probability. Then, we utilize self-identifiability to argue that if in some round $t$ of the main stage some suboptimal policy $\pi$ is chosen, there must exist some context-arm pair $(x,\pi(x))$ that is “under-explored”: appeared less than $\widetilde{O}(XK/\Gamma^{2}(f^{*})$ times. This step carefully harnesses the structure of the contextual bandit problem. Finally, we introduce a well-designed potential function (see Lemma E.7) that tracks the progress of learning over time. This function increases whenever a suboptimal policy is executed on an under-explored context-action pair, allowing us to bound the total number of times any suboptimal policy is executed. A key challenge is that while the second step guarantees the existence of an “under-explored” context-arm pair, it does not ensure that the context actually appears when the associated policy is chosen. We address this using a supermartingale argument and the fact that each context arrives with probability at least $p_{0}$ in each round. Combining these steps, we derive an upper bound on the expected number of times $\mathtt{Greedy}$ selects a suboptimal policy, and we bound the final expected regret via the regret decomposition lemma.
Proof Sketch for Theorem 4.3(b). As in the MAB case, we define event $E_{1}$ to ensure that the warm-up data misidentifies $f_{\mathtt{dec}}$ as the true reward function, and event $E_{2}$ that the empirical rewards of the decoy policy are tightly concentrated. The definitions are modified to account for contexts:
$E_{1}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(x,a)-f^{\dagger}(x,a)}<\Gamma(f^{\dagger})/2\quad\text{for each $x\in\mathcal{X}$ and arm $a\neq\pi^{\dagger}(x)$}\,\right\},$ (4.3)
$E_{2}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{t}(x,\pi^{\dagger}(x))-f^{*}(x,\pi^{\dagger}(x))}<\Gamma(f^{\dagger})/2\quad\text{for each $x\in\mathcal{X}$ and round $t>T_{0}$}\,\right\}.$ (4.4)
A decoy context-arm pair $(x,a)$ is one with $a=\pi_{\mathtt{dec}}(x)$. Event $E_{1}$ concerns the single warm-up sample for each non-decoy context-arm pair. Event $E_{2}$ asserts that the empirical rewards are concentrated for all decoy context-arm pairs (and all rounds throughout the main stage). Note that the two events are independent, as they concern non-overlapping sets of context-arm pairs.
$\mathtt{Greedy}$ always chooses the decoy arm when both events happen. We lower-bound $\Pr\left[\,E_{1}\cap E_{2}\,\right]$ by invoking independence and analyzing each event separately.
@@ 5 Interactive decision-making with arbitrary feedback
In this section, we consider the greedy algorithm in Decision-Making with Structured Observations ($\mathtt{DMSO}$). This is a general framework for sequential decision-making with a known structure, introduced in Foster et al. (2021), which allows for arbitrary feedback observed after each round, in addition to the reward. [^9] Bandit formulations with partial feedback that does not include the reward (known as partial monitoring, e.g.,  Bartók et al., 2014; Antos et al., 2013), are outside our scope. This feedback may be correlated with the respective rewards, which necessitates a modification of the greedy algorithm that looks beyond average/expected rewards.
Preliminaries. $\mathtt{DMSO}$ is defined as follows. Instead of “arms” and “contexts”, we have two new primitives: a policy set $\Pi$ and observation set $\mathcal{O}$. The interaction protocol is as follows: in each round $t=1,2,\,\ldots$, the algorithm selects a policy $\pi_{t}\in\Pi$, receives a reward $r_{t}\in\mathcal{R}\subset\mathbb{R}$, and observes an observation $o_{t}\in\mathcal{O}$. A model is a mapping from $\Pi$ to a distribution over ×$\mathbb{R}\times\mathcal{O}$. The reward-observation pair $(r_{t},o_{t})$ is an independent sample from distribution $M^{*}(\pi_{t})$, where $M^{*}$ is the true model. The problem structure is represented as a (known) model class $\mathcal{M}$ which contains $M^{*}$. We assume that $\Pi,\mathcal{M},\mathcal{R},\mathcal{O}$ are all finite.  [^10] Finiteness of the reward range $\mathcal{R}$ and observation set $\mathcal{O}$ is for ease of presentation. We can also handle infinite $\mathcal{R},\mathcal{O}$ if all outcome distributions $M(\pi)$ have a well-defined density, and 1 is stated in terms of these densities. To recap, the problem instance is a pair $(M^{*},\mathcal{M})$, where $\mathcal{M}$ is known but $M^{*}$ is not. This completes the definition of $\mathtt{DMSO}$.
@@
We’ll use the following notation. Let $f(\pi|M)$ be the expected reward for choosing policy $\pi$ under model $M$, and $f^{*}(\pi):=f(\pi|M^{*})$. A policy is called optimal (under model $M$) if it maximizes $f(\cdot\,|M)$, and suboptimal otherwise. Let $\pi^{*}$ be an optimal policy for $M^{*}$. The history $\mathcal{H}_{t}$ at round $t$ consists of $(\pi_{s},r_{s},o_{s})$ tuples for all rounds $s<t$. When two distributions $D,D^{\prime}$ are equal, we denote this as $D\stackrel{{\scriptstyle\text{d}}}{{=}}D^{\prime}$.
Modified Greedy. The modified greedy algorithm, called $\mathtt{GreedyMLE}$, uses maximum-likelihood estimation (MLE) to analyze reward-observation correlations. As before, the algorithm is initialized with $T_{0}\geq 1$ rounds of warm-up data, denoted $t\in[T_{0}]$. The warm-up data for a given round $t$ is a tuple $(\pi_{t},r_{t},o_{t})\in\Pi\times\mathbb{R}\times\mathcal{O}$, where the policy $\pi_{t}$ is chosen exogenously, and the $(r_{t},o_{t})$ pair is drawn independently from the corresponding distribution $M^{*}(\pi_{t})$. At each round $t>T_{0}$, the algorithm determines the highest-likelihood model given the history:
@@
where $\mathcal{L}(M|\mathcal{H}_{t})$ is the likelihood of model $M$ given history $\mathcal{H}_{t}$ and ties broken arbitrarily. [^12] Like in Section 2, we posit a “regression oracle” which implements the regression, and focus on statistical guarantees. Then the algorithm chooses the optimal policy given this model: $\pi_{t}\in\operatornamewithlimits{argmax}_{\pi\in\Pi}f(\pi|M_{t})$. For simplicity, we assume that the model class $\mathcal{M}$ guarantees no ties in this $\operatornamewithlimits{argmax}$.
@@
Here $\mathcal{L}(M|\mathcal{H}_{t})$ is an algorithm-independent notion of likelihood: the probability of seeing the reward-observation pairs in history $\mathcal{H}_{t}$ under model $M$, assuming that the policies in $\mathcal{H}_{t}$ have been chose in the respective rounds. In a formula,
@@
The equivalence of these definitions holds like in Claim 1:
@@
We define a suitable modification of function gap, termed model-gap, which takes the $\min$ over models and tracks the difference in reward-observation distributions (rather than expected rewards), expressed via KL-divergence (denoted $D_{\rm KL}$). Thus, the model-gap for model $M$ is as follows:
@@
We need an assumption on the ratios of probability masses:  [^13] Related (but incomparable) assumptions on mass/density ratios are common in the literature on online/offline reinforcement learning, (e.g.,  Munos and Szepesvári, 2008; Xie and Jiang, 2021; Zhan et al., 2022; Amortila et al., 2024).
@@
Thus, our characterization for $\mathtt{DMSO}$ is stated as follows:
@@
Fix a problem instance $(M^{*},\mathcal{M})$ of $\mathtt{DMSO}$ which satisfies 1. Let $\Gamma=\Gamma(M^{*},\mathcal{M})$ be the model-gap.
@@
Remark 5.4.
Again, $\mathtt{GreedyMLE}$ succeeds (i.e., achieves sublinear regret for any warm-up data) if and only if the problem instance is self-identifiable. Else, $\mathtt{Greedy}$ fails if the warm-up data consists of $N_{0}$ samples per each policy, as per Theorem 5.3(b), incurring linear regret like in Eq. (4.2).
We also provide a more efficient lower bound on $p_{\mathtt{dec}}$ in Theorem 5.3(b), where $B$ is replaced with a term that only concerns two models, $M_{\mathtt{dec}}$ and $M^{*}$. Letting $D_{\infty}$ be the Renyi divergence,
$p_{\mathtt{dec}}\geq e^{-O(C_{\mathtt{dec}}{N_{0}}|\Pi|)},\text{where }C_{\mathtt{dec}}=\max_{\pi\in\Pi}D_{\infty}\left(\,M_{\mathtt{dec}}(\pi)\,\|\,M^{*}(\pi)\,\right)\leq\log B.$ (5.3)
@@
We argue that suboptimal policies $\pi\in\Pi$ cannot be chosen “too often”. Indeed, fix one such policy $\pi$. Then with high probability the likelihood of any particular model $M\in\mathcal{M}_{\mathtt{opt}}(\pi)$ falls below that of $M^{*}$, and hence this model cannot be chosen again. Consequently, with high probability this policy cannot be chosen again, via a union bound over the models $M\in\mathcal{M}_{\mathtt{opt}}(\pi)$.  [^14] This last step is why $|\mathcal{M}|$ appears inside the $\log(\cdot)$ in the regret bound.
Part (b). We define independent events $E_{1}$ and $E_{2}$, resp., on the warm-up process and on all rounds when the decoy is chosen, so that $E_{1}\cap E_{2}$ guarantees that $\mathtt{GreedyMLE}$ gets forever stuck on the decoy. While this high-level plan is the same as before, its implementation is very different.
@@
First, we consider the “ghost process” for generating $\mathcal{H}_{\mathtt{warm}}$: in each round $t\leq T^{\prime}_{0}$, the chosen policy $\pi_{t}$ stays the same, but the outcome $(r_{t},o_{t})$ is generated according to the decoy model $M_{\mathtt{dec}}$. Under this process, each round raises the likelihood $\mathcal{L}_{t}(M_{\mathtt{dec}})$ more compared to any other model $M\in\mathcal{M}$. Specifically, write $\Delta\ell_{t}(M)=\Delta\ell_{t}(M\mid\mathcal{H}_{t})$ explicitly as a function of history $\mathcal{H}_{t}$, and let
@@
where the history is generated by the ghost process. Reusing Eq. (5.5) (with the decoy model $M_{\mathtt{dec}}$ now replacing the true model $M^{*}$), we obtain:
@@
happens with high probability when $\mathcal{H}_{\mathtt{warm}}$ comes from the ghost process. [^15] This argument invokes a concentration inequality, which in turn uses 1. Likewise, 1 is used for another application of concentration in the end of the proof sketch.
Since the ghost process and the true history $\mathcal{H}_{\mathtt{warm}}$ have bounded divergence, we argue that with some positive probability, event $E_{1}$ happens for the true $\mathcal{H}_{\mathtt{warm}}$, too. We use Renyi divergence $D_{\infty}$, see the full proof for details.

### v2

@@ 3 Characterization for structured bandits
We emphasize that self-identifiability and decoys are new notions, not reducible to structural notions from prior work, see Appendix B. It is easy to see that they are equivalent, in the following sense:
@@
In our characterization, the complexity of the problem instance $(f^{*},\mathcal{F})$ enters via its function-gap,
@@
Discussion. Thus, $\mathtt{Greedy}$ succeeds, in the sense of achieving sublinear regret for any warm-up data, if and only if the problem instance is self-identifiable. Else, $\mathtt{Greedy}$ fails for some warm-up data, incurring linear expected regret. Specifically, regret is $\operatornamewithlimits{\mathbb{E}}\left[\,R(t)\,\right]\geq(t-T_{0})\cdot p_{\mathtt{dec}}\cdot(f^{*}(a^{*})-f^{*}(a_{\mathtt{dec}}))$ for each round $t\in(T_{0},\infty)$, where $a^{*}$ is the best arm.
The correct perspective is that $\mathtt{Greedy}$ fails on every problem instance unless self-identifiability makes it intrinsically “easy”. Indeed, consider any bandit algorithm that avoids playing an arm once it is identified, with high confidence, as suboptimal and having a specific expected reward. This defines a mild yet fundamental non-degeneracy condition: a reasonable bandit algorithm should never take an action that provides neither new information (exploration) nor utility from existing information (exploitation), whether it prioritizes one or balances both. The class of algorithms satisfying this condition is broad—for instance, an algorithm may continue playing some arm $a$ indefinitely as long as the reward structure permits this arm to be optimal. However, under self-identifiability, any algorithm satisfying this condition achieves sublinear regret (see Section 8 for details).
The failure probability $p_{\mathtt{dec}}$ could be quite low. When there are multiple decoys $f_{\mathtt{dec}}\in\mathcal{F}$, we could pick one (in the analysis) which maximizes function-gap $\Gamma(f_{\mathtt{dec}})$. We present a more efficient analysis under stronger assumptions (which also applies to infinite function classes), see Section 6.
Proof Sketch for Theorem 3.3(a). We show that a suboptimal arm $a$ cannot be chosen more than $\widetilde{O}(K/\Gamma^{2}(f^{*}))$ times throughout the main stage. Indeed, suppose $a$ is chosen this many times by some round $t>T_{0}$. Then $\bar{r}_{t}(a)$, the empirical mean reward for $a$, is within $\Gamma(f^{*})/2$ of its true mean $f^{*}(a)$ with high probability, by a standard concentration inequality. This uniquely identifies $f^{*}(a)$ by definition of the function-gap, which in turn identifies $a$ as a suboptimal arm for any feasible reward function. Intuitively, this should imply that $a$ cannot be chosen again. Making this implication formal is non-trivial, requiring an additional argument invoking $\mathtt{MSE}_{t}(\cdot)$, as defined in (2.3).
@@
Proof Sketch for Theorem 3.3(b). To show that $\mathtt{Greedy}$ gets permanently trapped on the decoy arm despite reward randomness, we define two carefully-constructed events. The first ensures that the warm-up data causes $\mathtt{Greedy}$ to misidentify $f_{\mathtt{dec}}$ as the true reward function for all non-decoy arms:
@@
This concerns the single warm-up sample per non-decoy arm. The second event ensures that the empirical mean of the decoy arm $a_{\mathtt{dec}}$ remains close to $f^{*}(a_{\mathtt{dec}})$ for all rounds after the warm-up:
@@
Under $E_{1}\cap E_{2}$, $\mathtt{Greedy}$ always chooses the decoy arm. To lower-bound $\Pr\left[\,E_{1}\cap E_{2}\,\right]$, note that $E_{2},E_{1}$ are independent (as they concern, resp., $a_{\mathtt{dec}}$ and all other arms), analyze each event separately.
4 Characterization for structured contextual bandits ($\mathtt{StructuredCB}$)
The ideas from Section 3 need non-trivial modifications. The naive reduction to bandits — treating each contexts-to-arms mapping as a “super-arm” in $\mathtt{StructuredMAB}$— does not work because $\mathtt{Greedy}$ now observes contexts. Further, such reduction would replace the dependence on $K$ in Theorem 3.3 with the number of mappings, i.e., $K^{|\mathcal{X}|}$, whereas we effectively replace it with $K\cdot|\mathcal{X}|$.
Some notation: mappings from contexts to arms are commonly called policies. Let $\Pi$ denote the set of all policies. Expected reward of policy $\pi\in\Pi$ is $f^{*}(\pi):=\operatornamewithlimits{\mathbb{E}}_{x}\left[\,f^{*}(x,\pi(x))\,\right]$, where the expectation is over the fixed distribution of context arrivals. A policy $\pi$ is called optimal for reward function $f$ if it maximizes expected reward $f(\pi)$, and suboptimal otherwise. Let $\pi^{*}$ be the optimal policy for $f^{*}$. Note that $\pi(x)\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}f(x,\cdot)$ for each context $x$, which is unique by assumption. .
$\mathtt{Greedy}$ can be described in terms of policies: it chooses policy $\pi_{t}$ in each round $t$, before seeing the context $x_{t}$, and then chooses arm $a_{t}=\pi_{t}(x_{t})$. Here $\pi_{t}$ is the optimal policy for the $f_{t}$ from Eq. (2.1).
@@ 4 Characterization for structured contextual bandits ($\mathtt{StructuredCB}$)
Fix a problem instance $(f^{*},\mathcal{F})$ of $\mathtt{StructuredCB}$. Let $X=|\mathcal{X}|$.
@@
    Suppose the warm-up data consists of one sample for each context-arm pair. Assume $f^{*}$ has a decoy $f_{\mathtt{dec}}\in\mathcal{F}$, with decoy policy $\pi_{\mathtt{dec}}$. Then with some probability $p_{\mathtt{dec}}>0$, $\mathtt{Greedy}$ chooses $\pi_{\mathtt{dec}}$ in all rounds $t\in(T_{0},\infty)$. We have $p_{\mathtt{dec}}\geq X^{-O(KX/\Gamma^{2}(f_{\mathtt{dec}}))}$.
@@
$\mathtt{Greedy}$ succeeds (i.e., achieves sublinear regret for any warm-up data) if and only if the problem instance is self-identifiable. Else, $\mathtt{Greedy}$ fails for some warm-up data, with linear regret:
@@
New Proof Ideas. For Theorem 4.3(a), directly applying the proof techniques from the MAB case gives a regret bound linear in $|\Pi|=K^{X}$. Instead, we develop a non-trivial potential argument to achieve regret bound polynomial in $KX$. For Theorem 4.3(b), we give new definitions of events $E_{1},E_{2}$ extending (3.2), (3.3) by carefully accounting for contexts, and refine the deviation analysis to remove the dependence on $|\Pi|$. Proof sketches and full proofs are in Appendix D.
@@ 5 Interactive decision-making with arbitrary feedback
We consider Decision-Making with Structured Observations ($\mathtt{DMSO}$), a general framework for sequential decision-making with a known structure (Foster et al., 2021). It allows for arbitrary feedback observed after each round, along with the reward. [^7] Bandit formulations with partial feedback that does not include the reward (known as partial monitoring, e.g.,  Bartók et al., 2014; Antos et al., 2013), are outside our scope. This feedback may be correlated with the resp. rewards, necessitating a modification of $\mathtt{Greedy}$ which looks beyond average/expected rewards. The analysis becomes considerably more technical compared to $\mathtt{StructuredCB}$.
Preliminaries. $\mathtt{DMSO}$ is defined as follows. Instead of “arms” and “contexts”, we have two new primitives: a policy set $\Pi$ and observation set $\mathcal{O}$. The interaction protocol is as follows: in each round $t=1,2,\,\ldots$, the algorithm selects a policy $\pi_{t}\in\Pi$, receives a reward $r_{t}\in\mathcal{R}\subset\mathbb{R}$, and observes an observation $o_{t}\in\mathcal{O}$. A model is a mapping from $\Pi$ to a distribution over ×$\mathbb{R}\times\mathcal{O}$. The reward-observation pair $(r_{t},o_{t})$ is an independent sample from distribution $M^{*}(\pi_{t})$, where $M^{*}$ is the true model. The problem structure is represented as a (known) model class $\mathcal{M}$ which contains $M^{*}$. We assume that $\Pi,\mathcal{M},\mathcal{R},\mathcal{O}$ are all finite. [^8] Finiteness of $\mathcal{R},\mathcal{O}$ is for ease of presentation. We can also handle infinite $\mathcal{R},\mathcal{O}$ if all outcome distributions $M(\pi)$ have a well-defined density, and 1 is stated in terms of these densities. To recap, the problem instance is a pair $(M^{*},\mathcal{M})$, where $\mathcal{M}$ is known but $M^{*}$ is not. This completes the definition of $\mathtt{DMSO}$.
@@
We use some notation. Let $f(\pi|M)$ be the expected reward for choosing policy $\pi$ under model $M$, and $f^{*}(\pi):=f(\pi|M^{*})$. A policy is called optimal (under model $M$) if it maximizes $f(\cdot\,|M)$, and suboptimal otherwise. Let $\pi^{*}$ be an optimal policy for $M^{*}$. The history $\mathcal{H}_{t}$ at round $t$ consists of $(\pi_{s},r_{s},o_{s})$ tuples for all rounds $s<t$. $D\stackrel{{\scriptstyle\text{d}}}{{=}}D^{\prime}$ denotes that distributions $D,D^{\prime}$ are equal.
Modified Greedy. The modified greedy algorithm ($\mathtt{GreedyMLE}$) uses maximum-likelihood estimation (MLE) to analyze reward-observation correlations. As before, the algorithm is initialized with $T_{0}\geq 1$ rounds of warm-up data, denoted $t\in[T_{0}]$. Each round yields a tuple $(\pi_{t},r_{t},o_{t})\in\Pi\times\mathbb{R}\times\mathcal{O}$, where the policy $\pi_{t}$ is chosen exogenously, and the $(r_{t},o_{t})$ pair is drawn independently from the corresponding distribution $M^{*}(\pi_{t})$. At each round $t>T_{0}$, the algorithm determines
@@
the model with the highest likelihood $\mathcal{L}(M|\mathcal{H}_{t})$ given history $\mathcal{H}_{t}$ (with ties broken arbitrarily). [^10] As in Section 2, the regression is implemented via a “regression oracle’”; we focus on statistical guarantees. Then the algorithm chooses the optimal policy given this model: $\pi_{t}\in\operatornamewithlimits{argmax}_{\pi\in\Pi}f(\pi|M_{t})$. For simplicity, we assume that the model class $\mathcal{M}$ guarantees no ties in this $\operatornamewithlimits{argmax}$. Here $\mathcal{L}(M|\mathcal{H}_{t})$ is an algorithm-independent notion of likelihood: the probability of seeing the reward-observation pairs in history $\mathcal{H}_{t}$ under model $M$, if the policies in $\mathcal{H}_{t}$ were chosen in the resp. rounds. In a formula,
@@
We define model-gap, a modification of function gap which tracks the difference in reward-observation distributions (expressed via KL-divergence, denoted $D_{\rm KL}$). The model gap of model $M\in\mathcal{M}$ is
@@
Our characterization needs an assumption on the ratios of probability masses:  [^11] Related (but incomparable) assumptions on mass/density ratios are common in the literature on online/offline RL, (e.g.,  Munos and Szepesvári, 2008; Xie and Jiang, 2021; Zhan et al., 2022; Amortila et al., 2024).
@@
Fix an instance $(M^{*},\mathcal{M})$ of $\mathtt{DMSO}$ with 1 and model-gap $\Gamma=\Gamma(M^{*},\mathcal{M})$.
@@
$\mathtt{GreedyMLE}$ succeeds (i.e., achieves sublinear regret for any warm-up data) if and only if the problem instance is self-identifiable. Else, it fails for some warm-up data, with linear regret like in Eq. (4.2). We also provide a more efficient lower bound on $p_{\mathtt{dec}}$ in Theorem 5.3(b), replacing $B$ with a term that only concerns two relevant models, $M_{\mathtt{dec}},M^{*}$ (not all of $\mathcal{M}$). Letting $D_{\infty}$ be the Renyi divergence,
$p_{\mathtt{dec}}\geq e^{-O(C_{\mathtt{dec}}{N_{0}}|\Pi|)},\text{where }C_{\mathtt{dec}}={\textstyle\max_{\pi\in\Pi}}\;D_{\infty}\left(\,M_{\mathtt{dec}}(\pi)\,\|\,M^{*}(\pi)\,\right)\leq\log B.$ (5.3)
@@
We argue that suboptimal policies $\pi\in\Pi$ cannot be chosen “too often”. Indeed, fix one such policy $\pi$. Then with high probability (w.h.p.) the likelihood of any model $M\in\mathcal{M}_{\mathtt{opt}}(\pi)$ falls below that of $M^{*}$, so this model cannot be chosen again. So, w.h.p. this policy cannot be chosen again.  [^12] This last step takes a union bound over the models $M\in\mathcal{M}_{\mathtt{opt}}(\pi)$, hence $\log(\mathcal{M})$ in the regret bound.
Part (b). We define independent events $E_{1}$ and $E_{2}$, resp., on the warm-up process and on all rounds when the decoy is chosen, so that $E_{1}\cap E_{2}$ guarantees that $\mathtt{GreedyMLE}$ gets forever stuck on the decoy. While this high-level plan is the same as before, its implementation is far more challenging.
@@
First, we consider the “ghost process” ($\mathtt{ghost}$) for generating $\mathcal{H}_{\mathtt{warm}}$: in each round $t\leq T^{\prime}_{0}$, the chosen policy $\pi_{t}$ stays the same, but the outcome $(r_{t},o_{t})$ is generated according to the decoy model $M_{\mathtt{dec}}$. Under $\mathtt{ghost}$, each round raises the likelihood $\mathcal{L}_{t}(M_{\mathtt{dec}})$ more compared to any other model $M\in\mathcal{M}$. Namely, write $\Delta\ell_{t}(M)=\Delta\ell_{t}(M\mid\mathcal{H}_{t})$ explicitly as a function of history $\mathcal{H}_{t}$, and let
@@
where $\mathcal{H}_{t}$ comes from $\mathtt{ghost}$. Reusing Eq. (5.5) (with $M_{\mathtt{dec}}$ now replacing true model $M^{*}$), yields:
@@
happens w.h.p. when $\mathcal{H}_{\mathtt{warm}}$ comes from $\mathtt{ghost}$ [^13] This argument invokes a concentration inequality, which in turn uses 1. Likewise, 1 is used for another application of concentration in the end of the proof sketch. Since $\mathtt{ghost}$ and $\mathcal{H}_{\mathtt{warm}}$ have bounded Renyi divergence, we argue that with some positive probability, event $E_{1}$ happens under $\mathcal{H}_{\mathtt{warm}}$.